The area of a rectangle is represented by x2-4, and a pair of consecutive sides are measured x+4 and x-3. What is the area of the rectangle?
Answer by nerdybill(7384)

The area of a rectangle is represented by x^2-4, and a pair of consecutive sides are measured x+4 and x-3. What is the area of the rectangle?
.
Solve for x:
(x+4)(x-3) = x^2-4
FOIL left side:
x^2-3x+4x-12 = x^2-4
x^2+x-12 = x^2-4
x-12 = -4
x = 8
.
Area was given as:
x^2-4
plug in our value of x:
8^2-4
64-4
60 square units (answer)